olivia14:

What inequality will show the number of trucks used to transport the required raw material?

olivia14:

Brazyxkapalot:

200+15x=1500 I MIGHT B WRONG

olivia14:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Brazyxkapalot 200+15x=1500 I MIGHT B WRONG \(\color{#0cbb34}{\text{End of Quote}}\) okay thank u mamas lemme try it

Brazyxkapalot:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @olivia14 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Brazyxkapalot 200+15x=1500 I MIGHT B WRONG \(\color{#0cbb34}{\text{End of Quote}}\) okay thank u mamas lemme try it \(\color{#0cbb34}{\text{End of Quote}}\) Yw bae I think im wrong doe i think u need <,> with tha line undaneath